OpenStudy (anonymous):

HELP ME Solve 4(2x − 1) − 3 = 9(x − 4). −29 32 13 29

OpenStudy (usukidoll):

distribute 4(2x − 1) and 9(x − 4). first

OpenStudy (anonymous):

okay 6x-4*1???? and 9x*9*4

OpenStudy (anonymous):

i meant - in the middle of the second one

OpenStudy (usukidoll):

o-o distribute means to multiply what is 4 times 2x and what is 4 x 1 ?

OpenStudy (anonymous):

8x and 4

OpenStudy (usukidoll):

and 4 times 1 similarly what is 9 times x and 9 x 4 9(x − 4)

OpenStudy (usukidoll):

ack ignore first sentence...

OpenStudy (anonymous):

lol 9x and 36

OpenStudy (usukidoll):

4(2x − 1) − 3 = 9(x − 4). 8x-4-3=9x-36

OpenStudy (usukidoll):

so we need to clean this up a bit ... what is -4-3?

OpenStudy (anonymous):

-7

OpenStudy (usukidoll):

yay 8x-4-3=9x-36 8x-7=9x-36 now I need to have all x's to the left... what do we need to do next?

OpenStudy (anonymous):

add 36??

OpenStudy (usukidoll):

umm yeah but we're not there yet... let's have all x's to the left first .

OpenStudy (usukidoll):

which terms have x in them?

OpenStudy (usukidoll):

8x-7=9x-36

OpenStudy (anonymous):

8x and 9x

OpenStudy (usukidoll):

yes... and we need all variables (that is the letters) to the left since 8x is already on the left... what do we need to do to have 9x on the left?

OpenStudy (anonymous):

add?

OpenStudy (usukidoll):

ummm no we need to get rid of the 9x on the right so we do what to both sides?

OpenStudy (anonymous):

divviidddee?? No, subtracct.. No, divide.

OpenStudy (usukidoll):

we don't do division...we do subtraction

OpenStudy (usukidoll):

so we subtract 9x on both sides 8x-7-9x=9x-9x-36 the 9x on the right should go to 0 then we rewrite this

OpenStudy (usukidoll):

8x-9x-7=-36

OpenStudy (usukidoll):

what is 8x-9x ?

OpenStudy (anonymous):

-1x

OpenStudy (mathstudent55):

@Marianakay456 On the right side you have 9x - 36 Think of it as + 9x - 36 The 9x is being added to the right side, the 36 is being subtracted from the right side. Since you want to get rid of the +9x from the right side, you subtract 9x from both sides.

OpenStudy (usukidoll):

yeah -x-7=-36 now I need all numbers to go on the right... what do we need to do to the 7?

OpenStudy (anonymous):

add?

OpenStudy (usukidoll):

yes we add 7 to both sides -x-7+7=-36+7 -x = -36+7

OpenStudy (usukidoll):

what is -36+7 ?

OpenStudy (anonymous):

-29

OpenStudy (usukidoll):

-x = -29 yes... there is one more step though....what is it?

OpenStudy (anonymous):

do we make it positive? lol idk

OpenStudy (usukidoll):

yeah we make it positive by dividing ???????? because I want x = ? to be my final answer

OpenStudy (anonymous):

-29?

OpenStudy (anonymous):

so the answer is +29 or just 29

OpenStudy (usukidoll):

-x = -29 how to make this positive.... we use division .. divide by ??? to have x = ?

OpenStudy (anonymous):

is D correct? 29 or +29?

OpenStudy (mathstudent55):

-x means -1 * x divide both sides by ...

OpenStudy (anonymous):

-1

OpenStudy (usukidoll):

please understand the last step before choosing the answer -x=-29 there is a step missing.

OpenStudy (usukidoll):

division is required... divide ____________ on both sides

OpenStudy (anonymous):

we need to divide by -1

OpenStudy (usukidoll):

yes on both sides

OpenStudy (anonymous):

OKK. we end up with DUNDUNDUN my answer, D

OpenStudy (usukidoll):

yeah

OpenStudy (anonymous):

thanks